OpenStudy (anonymous):

evaluate. (b^3)^2 b^5 ya?

OpenStudy (saifoo.khan):

Ya

Parth (parthkohli):

Nope!

OpenStudy (saifoo.khan):

Wait! NO!

mathslover (mathslover):

\(\large{(a^b)^c=a^{bc}}\)

OpenStudy (anonymous):

No..

Parth (parthkohli):

\( \color{Black}{\Rightarrow n^3 \times n^3 = n^{3 + 3}}\)

mathslover (mathslover):

hence it will be : \(\large{(b^3)^2=b^6}\)

OpenStudy (saifoo.khan):